You are given vs = A1.v2.cos (100t + B1) Vc = A2 · cos (100t + B2) Find VR = Az · cos (100t + B3) with – 180° 5 B3 = 180° + VPlease explain how you added the phasors together.
Answer #1

Sautions Given data, Vs = A. . 12. cos (1oof AB) Ne = A2. Cos Eloot tB] A,=2V, Bp=25°, A2 = 20 , By=-28 - 1 15 = 2v1 cos (loVR = 2 cos Cloot + 802) - VR = A3 cos (lost it B3) By comparing go & ego A-2 volts Baso degrees